November 1997  ICTCM X (Chicago)  6

More Challenging Problems

• multi-point boundary conditions y′′ + y = 0, y(π/4) + y(π/2) = 1, y′(π/4) + y′(π/2) = −1

• nonlinear boundary conditions y′′ + y = 0, y(0) = 2, y(π/2)^{2} + y′(π) = 1

• nonlinear differential equation y′′ + y^{2} = 0, y(0) = 2, y(π/2) = 0